OpenStudy (anonymous):

Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)).

OpenStudy (zale101):

\[f(x^2-7x-9)\] since g(-1) that means g of -1. So, replace the x values with -1 and then compose it with the function of f.

OpenStudy (zale101):

\[g=(-1)^2-7(-1)-9\]

OpenStudy (anonymous):

These are the answer choices. I don't get how that works? -21 -3 3 21

OpenStudy (jdoe0001):

\(\bf f({\color{brown}{ x}})=x-2\qquad {\color{brown}{ g(x)}}=x^2 - 7x - 9 \\ \quad \\ f(\quad {\color{brown}{ g(x)}}\quad )={\color{brown}{ g(x)}} -2\implies f(\quad {\color{brown}{ g(x)}}\quad )={\color{brown}{ (x^2 - 7x - 9)}} -2 \\ \quad \\ f(\quad {\color{brown}{ g({\color{blue}{ -1}})}}\quad )={\color{brown}{ (({\color{blue}{ -1}})^2 - 7({\color{blue}{ -1}}) - 9)}} -2\)